how do I solve:
3x/5 = -6
y/4 + 2 = 7
16 = -2x/3
Find f(1) for f(x) = 2x3 - 3x2 + x – 21

Answer by RAY100(1637)

3x/5 = -6,,,,mult by 5
3x=-30,,,,,,divide by 3
x=-10
.
.
y/4 +2 =7,,,,,subst 2
y/4 =5,,,,,,,,,mult by 4
y=20
.
.16 = -2x/3,,,,,,,,mult by 3
48 =-2x,,,,,,,,,,,divide by (-2)
-24 =x
.
.
f(x) =2x^3 -3x^2 +x -21
f(1) = 2(1)^3 -3(x)^2 +(1) -21
f(1)= 2-3+1-21 = -21
.